OpenStudy (anonymous):

write an equation in slope intercept form of the line that passes through the given point and is perpendicular to the graph of the given equation. (0,0) y=-3x+2 (5,0) y+1=2(x-3)

OpenStudy (anonymous):

@Data_LG2

OpenStudy (anonymous):

so now, they are looking for the "perpendicular line" remember what i said before about their slope? the slope of a perpendicular line is the \(\sf negative\ reciprocal\) of the given line. so the first one, what is the slope?

OpenStudy (anonymous):

-3x

OpenStudy (anonymous):

right, just take the coefficient. the slope is -3 now what is the negative reciprocal of -3?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

close ! how about it's reciprocal? we have 3 and underneath it, we have the imaginary 1 so to get the reciprocal, we have to flip them look here: (give me a sec my drawing tool is not working)

OpenStudy (anonymous):

1/3

OpenStudy (anonymous):

yes right, (lol i don't need the drawing tool then xD ) the slope of the required line is 1/3 so now we have \(\sf y=\frac{1}{3}x+b\) like before, we have to solve for \(\sf b\)

OpenStudy (anonymous):

b=0

OpenStudy (anonymous):

@Catlover5925

OpenStudy (anonymous):

can you help me with the second one plz :3

OpenStudy (catlover5925):

i can try, what is the question??

OpenStudy (anonymous):

^^same question at the top just the second equation now

OpenStudy (catlover5925):

ummm i am not sure @Data_LG2 said he would be right back...sorry

OpenStudy (anonymous):

its ok.

OpenStudy (catlover5925):

@mathmath333

OpenStudy (mathmath333):

(0,0) y=-3x+2 \(\large\tt \begin{align} \color{black}{ y=-3x+2\\~\\ }\end{align}\) the line passing through (0,0) and perpendicular with y=-3x+2 have its slope as negative inverse of the line y=-3x+2 so \(m=\dfrac{1}{3}\) \(\large\tt \begin{align} \color{black}{( y-y_1)=m(x-x_1)\\~\\ ( y-0)=\dfrac{1}{3}(x-0)\\~\\ y=\dfrac{1}{3}x\\~\\ }\end{align}\) https://www.desmos.com/calculator/kt7mmqvgif